OpenStudy (anonymous):

someone please factor this completely 6(x+5)^2-5(x+5)+1

OpenStudy (mertsj):

[(3(x+5)-1][2(x+5)-1]

OpenStudy (anonymous):

Can you show me step by step how you did it please?

OpenStudy (mertsj):

Could you factor this: 6a^2-5a+1

OpenStudy (anonymous):

ya i got (1 + -3a)(1 + -2a) is that right?

OpenStudy (mertsj):

Yes or more typically (3a-1)(2a-1)

OpenStudy (anonymous):

oh ok. so how did you get that expression i asked?

OpenStudy (anonymous):

would you mind checking my other factoring questions i answered because I am not sure if its right?

OpenStudy (mertsj):

Now, notice your problem is exactly the same as the one you just factored. Just replace each a with (x+5)

OpenStudy (anonymous):

replace it where?

OpenStudy (mertsj):

6 a^2 -5 a +1 = (3 a - 1)(2 a - 1) 6(x+5)^2-5(x+5) +1 = (3(x+5)-1)(2(x+5) -1 )

OpenStudy (mertsj):

Do you see how the two problems are exactly the same?

OpenStudy (anonymous):

oh yes i do